Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point? North and east can be considered the directions of the y- and x-axis respectively.
Answer:
Show work in this space.

Answer by checkley71(8403)

35^2+65^2=D^2 OR 1225+4225=D^2 OR 5450=D^2 OR D=73.824115 KILOMETERS FROM THE STARTING POINT.